OpenStudy (anonymous):

simplify: sin 2x cos 2x using derivatives.

OpenStudy (anonymous):

That makes no sense. Can you please reword that?

OpenStudy (anonymous):

do you want us to help you take the derivative of that function?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

i would clean it up a little first using the double angle formula: \[\sin(2x)\cos(2x) = \frac{1}{2}(2\sin(2x)\cos(2x)) = \frac{1}{2}\sin(4x)\]

OpenStudy (anonymous):

Now taking the derivative is just an application of the chain rule, instead of chain and product rule. \[\frac{d}{dx}\left(\frac{1}{2}\sin(4x)\right) = \frac{1}{2}\cos(4x)*4 = 2\cos(4x)\]

OpenStudy (anonymous):

can i ask how come a one half existed in the simplification?

OpenStudy (anonymous):

i multiplied by a clever form of "one": \[\sin(2x)\cos(2x) = 1*\sin(2x)\cos(2x) = \frac{2}{2}\sin(2x)\cos(2x) = \]\[\frac{1}{2}(2\sin(2x)\cos(2x))\]